# Bootstrap 3D Reconstructed Scenes from 3D Gaussian Splatting

### v1

@@ 4.2 Results Analysis
Our improvement is evident not only in the metrics. As previously mentioned, we also addressed issues such as artifacts that occur when changing the camera zoom or rendering views that significantly diverge from the training sets [31] to some extent. Bootstrapping mitigates artifacts observed in the original 3D-GS by generating pseudo-details as shown in Fig 5. By the conclusion of our third-stage experiments, we observed that artifacts occurring in specific views could indeed be alleviated while maintaining strong performance metrics. However, there remains considerable potential for further improvement. Given that our technique has already achieved superior results, our fourth-stage experiments will primarily focus on further alleviating these artifacts. Details of our fourth-stage experiments are illustrated in our appendix. From Fig 1 and 6, it is evident that our fourth-stage results show clear improvements over both the original 3D-GS and our third-stage outcomes. Distortions observed in the undersampled view and the jumbled mass under the table have been substantially modified. After completing our fourth-stage experiments on the Mip-NeRF 360 datasets [83], we observed that the application of the upscale diffusion model is particularly effective in open scenes that encompass many unseen views (maybe because we do not finetune it). For most indoor scenes, the introduction of significant noise through the upscale diffusion model tends to negatively impact the performance metrics substantially, as detailed in our appendix.
@@
Thanks to the 3D-GS’s fast training and adaptation to reconstruction, our technique can have an opportunity to make use of its potential. The central increase in time consumption for our technique stems from the diffusion process introduced by bootstrapping. On the other hand, the multi-view rasterization, as required in Eq. 8, has minimal impact on the overall time required. Our time consumption comparisons are exhibited in Tab 2. Generally, the increase in diffusion time steps, which is related to the broken strength $s_{r}$, along with the use of the upscale diffusion model in our fourth-stage experiments, significantly contributes to additional training time.
@@
We also made comparisons of rendering details across our four-stage experiments. During the initial two stages, if we evaluate solely based on the generation quality without considering the ground truth, there appears to be minimal variance in their quality when observed with the naked eye. After fine-tuning for each specific scene, we observed significant progress in context-specific details. For instance, in Fig 4 Room, consider the enclosed camera above the door. In our first-stage experiment, the camera was only represented in a basic form resembling a camera. However, in our second-stage experiment, the camera displayed curves similar to those in the ground truth. Then in our third-stage experiments, although the details were altered and refined due to increased DDIM sampling steps, as seen in Fig 5, these alterations were unable to fully compensate for the substantial deficiencies present in the scene. Furthermore, these generated details tend to occur at low frequencies, which is closely related to our loss term. Since our approach emphasizes enhancing details that harmonize rather than conflict with each other, the high-frequency details generated by the diffusion model tend to exhibit variability. Ultimately, while our fourth-stage experiments were capable of significantly modifying the unseen content, these changes did not necessarily translate into improved metrics. This is because the fine-grained details generated might not always accurately reflect the actual ground truth details.
@@ 4.3 Ablation Studies
We focus our ablation studies mainly on the configuration of the diffusion model, specifically on how we regenerate the bootstrapping images. We have evaluated various hyperparameters of the diffusion model, with some outcomes outlined in Tab 3. Our results indicate that only the broken strength $s_{r}$ and the loss weight $\lambda_{\text{boot}}$ are significant. Common hyperparameters that typically influence the outcomes in traditional text-to-image generation, such as text prompt, text prompt scale, and eta as described in [30], have minimal impact in our experimental setup. As detailed in Sec. 4.1, we apply our bootstrapping technique only during specific intervals. We experimented with conducting consecutive bootstrapping after the 6,000th iteration but discovered that this approach led to instability. Continuous bootstrapping, when overused, tends to generate scenes that are completely out of context. Even when finetuning the diffusion model specifically for each scene, its instability remains a significant issue. The highest results achieved through consecutive bootstrapping in the second-stage setting occasionally surpass those of our best records. However, achieving these superior outcomes is quite challenging, and the training time required is approximately 2.5 times longer compared to our third-stage experiments. This significant increase in time and effort highlights the inefficiency of continuous bootstrapping despite its potential for achieving slightly better performance metrics. Notice that our other basic ablation studies are all tested on our first-stage experiment’s configuration.
@@ 4.4 Discussions and Limitations
Through our experiments, we identified that the main challenge in implementing our approach is the time consumption. Achieving desirable results often necessitates extended periods in the diffusion process across different experimental stages. Specifically, while we can achieve excellent outcomes in our second-stage experiments, training the 3D-GS model requires hours. However, given that our technique is designed as a plug-and-play solution, we anticipate it will evolve and improve as advancements are made in this field. Further details are discussed in the supplementary material.
@@ 8 Experiments and Results
Here we list the error metrics used in our evaluation in Sec.4 across all considered methods and scenes, as shown in Tab. 4-9.

### v2

@@ 4.2 Results Analysis
Our improvement is evident not only in the metrics. As previously mentioned, we also addressed issues such as artifacts that occur when changing the camera zoom or rendering views that significantly diverge from the training sets [31] to some extent. Bootstrapping mitigates artifacts observed in the original 3D-GS by generating pseudo-details as shown in Fig 6. By the conclusion of our third-stage experiments, we observed that artifacts occurring in specific views could indeed be alleviated while maintaining strong performance metrics. However, there remains considerable potential for further improvement. Given that our technique has already achieved superior results, our fourth-stage experiments will primarily focus on further alleviating these artifacts. Details of our fourth-stage experiments are illustrated in our appendix. From Fig 1 and 7, it is evident that our fourth-stage results show clear improvements over both the original 3D-GS and our third-stage outcomes. Distortions observed in the undersampled view and the jumbled mass under the table have been substantially modified. After completing our fourth-stage experiments on the Mip-NeRF 360 datasets [83], we observed that the application of the upscale diffusion model is particularly effective in open scenes that encompass many unseen views (maybe because we do not finetune it). For most indoor scenes, the introduction of significant noise through the upscale diffusion model tends to negatively impact the performance metrics substantially, as detailed in our appendix.
@@
Thanks to the 3D-GS’s fast training and adaptation to reconstruction, our technique can have an opportunity to make use of its potential. The central increase in time consumption for our technique stems from the diffusion process introduced by bootstrapping. On the other hand, the multi-view rasterization, as required in Eq. 9, has minimal impact on the overall time required. Our time consumption comparisons are exhibited in Tab 2. Generally, the increase in diffusion time steps, which is related to the broken strength $s_{r}$, significantly contributes to additional training time.
@@
We also made comparisons of rendering details across our four-stage experiments. During the initial two stages, if we evaluate solely based on the generation quality without considering the metrics, there appears to be minor variance in their quality. After fine-tuning for each specific scene, we observed significant progress in context-specific details. For instance, in Fig 5 Room, consider the camera object (the middle one in zoomed-in details) above the door. In our first-stage experiment, the camera was only represented in a basic form resembling a camera. However, in our second-stage experiment, the camera displayed curves similar to those in the ground truth. Then in our third-stage experiments, although the details were altered and refined due to increased DDIM sampling steps, as seen in Fig 6, these alterations were unable to fully compensate for the substantial deficiencies present in the scene. Furthermore, these generated details tend to occur at low frequencies, which is closely related to our loss term. Since our approach emphasizes enhancing details that harmonize rather than conflict with each other, the high-frequency details generated by the diffusion model tend to exhibit variability. Ultimately, while our fourth-stage experiments were capable of significantly modifying the unseen content, these changes did not necessarily translate into improved metrics. This is because the fine-grained details generated might not always accurately reflect the actual ground truth details.
@@ 4.3 Ablation Studies
Our ablation studies are mainly tested on the configuration of the diffusion model, specifically on how we regenerate the bootstrapping images. We have evaluated various hyperparameters of the diffusion model, with some outcomes outlined in Tab 3. Our results indicate that only the broken strength $s_{r}$ and the loss weight $\lambda_{\text{boot}}$ are significant. Common hyperparameters that typically influence the outcomes in the traditional text-to-image generation, such as text prompt, text prompt scale, and eta as described in [30], have minimal impact in our experimental setup. As detailed in Sec. 4.1, we apply our bootstrapping technique only during specific intervals. We have also conducted consecutive bootstrapping with every bootstrapping interval after the 6,000th iteration but discovered that this approach led to instability. Continuous bootstrapping, when overused, tends to generate scenes that are completely out of context. Even when finetuning the diffusion model specifically for each scene, its instability remains a significant issue. The highest results achieved through consecutive bootstrapping in the second-stage setting occasionally surpass those of our best records. However, achieving these superior outcomes is quite challenging, and the training time required is approximately 2.5 times compared to the original second-stage experiments. This significant increase in time and effort highlights the inefficiency of consecutive bootstrapping despite its potential for achieving slightly better performance metrics. Notice that our other basic ablation studies are all tested on our first-stage experiment’s configuration.
@@ 4.4 Discussions and Limitations
Through our experiments, we identified that the main challenge in implementing our approach is the time consumption. Achieving desirable results often necessitates extended periods in the diffusion process across different experimental stages. Specifically, while we can achieve excellent outcomes in our second-stage experiments, training the 3D-GS model requires hours. In addition, we extended our approach to incorporate multi-scale 3D datasets [51]. While this extension yielded promising results in comparison to the original 3D-GS, it also revealed notable limitations, as outlined in our appendix. However, given that our technique is designed as a plug-and-play solution, we anticipate it will evolve and improve as advancements are made in this field. Further details are discussed in the supplementary material.
@@ 7.3 Implementation of Upscale Diffusion Model
7.4 Implementation of Multi-scale 3D Datasets
Multi-scale datasets [51] pose substantial challenges to the original 3D-GS since this model primarily replicates the visible expressions in training images without adequately considering the underlying structure. Consequently, distortions frequently occur in a well-trained 3D-GS model when rendering from a close perspective. In our approach, we utilize the evaluation perspective as bootstrapping views and simultaneously expand these views. Initially, we gather all camera information pertaining to the evaluation views, followed by a Consecutive expansion of these views. During training, we do not integrate these bootstrapped, evaluation-close views with their corresponding far-training views. Instead, we treat these bootstrapped views directly as training inputs. For each evaluation-close view, we employ multiple consecutive renderings for training, following the same methodology as outlined in our loss term, Eq. 9. We also constrain this loss term with a scale factor of 0.1 to maintain training stability. After experiments, we discovered that only the finetuned SD2.1v model is suitable for use with these datasets [51], so we also finetuned the SD2.1v model for each scene. In the finetuning phase, we trained two 3D-GS models: one utilizing the full datasets without segregating for evaluation, and the other maintaining the standard evaluation setup. We bootstrapped these models and simultaneously rendered both evaluation and bootstrapping views on these two models. After that, the renderings from the first model were considered as the ground truth to train the renderings of the second model. Although we achieved significant advancements over the original 3D-GS model, our training strategy involving the bootstrapping of renderings could be considered somewhat deceptive, as our diffusion model has already been exposed to the evaluation datasets. As a result, we have opted not to include this part in our main paper. However, this approach does demonstrate the effectiveness of our bootstrapping technique, validating its potential under certain conditions.
@@ 8 Experiments and Results
Here we list the error metrics used in our evaluation in Sec.4 across all considered methods and scenes, as shown in Tab. 4-10.
@@
Table 10: PSNR scores for BungeeNeRF [51].
Method Scenes
Amsterdam Bilbao Pompidou Quebec Rome Hollywood
3D-GS [1] 25.74 26.35 21.20 28.79 23.54 23.25
Scaffold-GS [35] 27.10 27.66 25.34 30.51 26.50 24.97
Ours 26.85 28.13 24.79 31.15 26.77 25.02